OpenStudy (campayne):

Write the equation of the line through (-1, 0) and (1, 2) in slope-intercept form.

OpenStudy (imnotarobot5):

Is it multiple choice?

OpenStudy (campayne):

A) y = x + 1 B) y = m(x - x1) + 1 C) y - y1= m(x - x1) + 1 D) y - y1= (x - x1) + 1

OpenStudy (imnotarobot5):

Alright, do you know how to get the slope?

OpenStudy (campayne):

y= mx + b?

OpenStudy (imnotarobot5):

Nah like the slope of the line. So the equation y = mx +b will be your equation for a specific line

OpenStudy (imnotarobot5):

but the slope is how steep that line is, its a part of the equation but not the whole equation

OpenStudy (campayne):

well.. im kinda lost

OpenStudy (imnotarobot5):

Haha, that's okay. Do you want me to kind of break down the subject a little? Maybe that'll help you understand the rest of the equations you have to do a little better

OpenStudy (campayne):

that could help

OpenStudy (imnotarobot5):

Alright, so what we're finding is the equation of a line

OpenStudy (imnotarobot5):



OpenStudy (imnotarobot5):

SO lets say thats those two points, you can base a line off them, and that line would constitute an equation to go with it

OpenStudy (imnotarobot5):